OpenStudy (anonymous):

Jefferson Park sits on one square city block 300 ft on each side sidewalks across the park join opposite corners. To the nearest foot, how long is each diagonal sidewalk?

OpenStudy (anonymous):

Diagonals of squares are sqrt(2) times the side length

OpenStudy (anonymous):

that doesnt help i need a like written explanation

OpenStudy (anonymous):

i got 424 is that right?

OpenStudy (anonymous):

Just so.

OpenStudy (anonymous):

Square has sides S Diagonal is hypotenuse of right triangle of sides S and S Length of hypotenuse squared = h^2 = S^2 + S^2 = 2 S^2 h = sqrt(2) S